compute its eigen value and vector

oto 19. -2
Answer #1

Cet -1 eigen value (2-d) [trd)(1-d) -o] (2-d) Eld+1) (d-1)] = 0 +o - 0 -0 +0=O d= 2,1,- eigen vector (A- 21) V, = 0 (::)G) -(A+I) V3 = 0 34 =0 ç=0 4-2C3 =0 -4 7213 =0 3 ) (3=0 Take V3=